OpenStudy (anonymous):

A, B and C can do a piece of work in 20, 30 and 60 days respectively. In how many days can A do the work if he is assisted by B and C on every third day?

OpenStudy (cwrw238):

when working together 1/20 + 1/30 +1/60 = 1/x where x = number of days to do the job

OpenStudy (anonymous):

still cant find the answer. :(

OpenStudy (cwrw238):

its tricky this one A works for 2 days then 3 of them work for one and so on...

OpenStudy (anonymous):

appreciate the effort miss. thank you :D but i really have to answer this. -sniffs-

OpenStudy (cwrw238):

so A does 1/10 of job in 2 days then - ok - no problem

OpenStudy (anonymous):

yes.

OpenStudy (cwrw238):

x = 10 so in 3 days 1/10 + 1/10 = 1/5 of job is done all you have to do now is add up

OpenStudy (anonymous):

thanks miss. imma do that.

OpenStudy (cwrw238):

yw